OpenStudy (anonymous):

Prime Factorization 35^3 *65*9^13 where do I start?

zepdrix (zepdrix):

Ummm I think you want to break down the bases into their prime factorization, then combine powers and such. So for example the 35 would be.. \[35^3=(5*7)^3=5^3*7^3\] 5 and 7 are prime, so that number is finished, yay! Can you figure out the primes of 65 and 9? :)

OpenStudy (anonymous):

5,13 for 65 3,3 for 9

zepdrix (zepdrix):

\[(5^3*7^3)(5^1*13^1)(3^{15}*3^{15})\] Ok now combine terms with similar bases! Remember when you multiply terms with similar bases, you ADD the exponents

OpenStudy (anonymous):

5*13*5^3*7^3*13^26

OpenStudy (anonymous):

how do you use exponents?

zepdrix (zepdrix):

\[3^{15}*3^{15}=3^{15+15}=3^{30}\] When you multiply terms with the same base (3 is our base), You ADD the exponents, as I have done here.